OpenStudy (anonymous):

solve a-5 5 ____ = _ a+4 4 a = 0 a = -1 a = -9 a = -40

OpenStudy (anonymous):

a-5=5 a+4=4 is it like this?

OpenStudy (anonymous):

if yes then for first part: a-5=5 a=5+5 a=10 for second one a+4=4 a=4-4 a=0